# LEADRE: Multi-Faceted Knowledge Enhanced LLM Empowered Display Advertisement Recommender System

### v1

@@ 3.3. Trie-Tree Construction
In the process of ad indexing, each ad $a$ is represented by a list of semantic tokens denoted as $\hat{\mathbf{c}}_{a}=[c_{a}^{1},c_{a}^{2},\dots,c_{a}^{M}]$. To handle potential collisions where different ads may map to the same S-IDs, we introduce an additional code to distinguish them (Rajput et al., 2024). As a result, the S-IDs for each ad are updated to $\mathbf{c}_{a}=[c_{a}^{1},c_{a}^{2},\dots,c_{a}^{L},c_{a}^{(M+1)}]$, where $c_{a}^{(M+1)}$ represents the additional code used to disambiguate ads with identical list of S-IDs. To distinguish different level semantic tokens, we use ¡a_ , b_, c_, $\cdots$ ¿ as prefix of different S-IDs level.
Next, we represent all ads within a trie-tree, where each path from the root node to a leaf node corresponds to the S-IDs of a specific ad. The Trie-Tree Construction Algorithm (detailed in Appendix A) iteratively adds each semantic token of an ad’s S-IDs as a child node, creates new nodes when necessary and marks the end of each complete ad’s S-IDs. This hierarchical structure enables efficient organization of ads, allowing for quick retrieval and prefix-based searches based on S-IDs.
@@ 3.4. Constrained Decoding
Different from conventional LLM-based chatting bot, generative retrieval for advertising system requires that the generated ads are valid ads in the predefined ads set, and should be a list of diverse ads instead of only one. To meet these requirements, we employ constrained decoding with the trie-tree (Hokamp and Liu, 2017; Hu et al., 2019; Post and Vilar, 2018). The Constrained Decoding Algorithm, detailed in Appendix A, iteratively expands candidate S-IDs lists through layers of the trie-tree, selecting the top $B$ lists at each step based on their scores, until either all layers are processed or the maximum length is reached. This approach ensures that the final generated ads are both high-probability according to the LLMs and valid within the predefined ad set.
@@ 4.1. Overview
LEADRE framework consists of three novel modules: (1) Intent-Aware Prompt Engineering designs intent-aware ¡Prompt, Response¿ pairs which accommodate user profile and behavior sequences, and are used as fine-tuning corpus for LLMs to learn users’ interests. (2) Advertising-Specific Knowledge Alignment incorporates auxiliary fine-tuning tasks and Direct Preference Optimization (DPO) to bridge the knowledge gap within LLMs and advertising system. (3) Efficient System Deployment integrates both latency tolerant service system and latency sensitive service system into LLMs deployment, ensuring scalability and real-time performance.
@@ 4.2.1. Prompt Components
Given their extensive training on vast amounts of web-crawled corpus, LLMs are capable of understanding and operating with content from various domains, as long as they are encoded as human understandable text in natural languages. We model user commercial intent by combining both long-term interests (derived from the user’s profile and historical behaviors) and short-term interests (based on recent behaviors). The prompt is assembled with the following key components:
@@
The following is an instruction describing a task. Please give a response to complete this request appropriately.
[Instruction]:¡Learnable Token¿Assuming you are an ad recommender system. 22-year-old male, resident in Haidian, Beijing, with a bachelor’s degree, working in Internet industry, with a medium consumption level. The categories that have been frequently interacted recently are (format: category^interaction times): emotion^115 times; entertainment^28 times; mental health^8 times; education^6 times; finance^6 times; The most recent interaction behavior sequence details (format: time^behavior type^title) are 32 days ago^play short video^reality; 31 days ago^click on ad^¡a_51, b_10, c_67, d_93, e_0¿; 27 days ago^play short video^shuttlecock; 25 days ago^Play short video^Emotional/psychological age test;22 days ago^Conversion ad^¡a_243, b_136, c_23, d_245, e_0¿;19 days ago^Click ad^¡a_164, b_243, c_38, d_88, e_0¿;16 days ago^Conversion ad^¡a_51, b_10, c_67, d_93, e_0¿. what ad will the user be interested in next time?
@@ 4.3.1. Semantic Alignment
Apart from the main task, we introduce auxiliary tasks to align the ad system with the language model and enhance the model’s advertising knowledge(Zheng et al., 2024). The auxiliary tasks consist of two components: an explicit alignment task and an implicit alignment task. These tasks not only improve the language model’s understanding of ads, but also contribute to better performance in the main task of the next ad generation.
The explicit alignment task is designed to predict the S-IDs of an ad based on its detailed description. This task directly aligns the language model with the ad system by making the LLMs understand how textual descriptions map to the corresponding ad S-IDs. A sample prompt-response pair for this task is shown below:
@@
In this example, LLM is provided with a detailed textual description of an ad, including the product’s name, type, brand, and category. The model’s task is to map this description to the corresponding S-IDs, which are internal identifiers used by the ad system.
The implicit alignment task builds on the Next Ad Generation mechanism described in the main task but replaces the S-IDs in the user’s ad interaction history with the ad descriptions. Instead of predicting the next ad based on a sequence of S-IDs (as in the main task), the model now predicts the next ad using textual descriptions of the ads. This subtle modification enhances the LLMs’ ability to associate ad descriptions with user behaviors, further aligning the model with the advertising system in a more implicit manner. By incorporating these auxiliary tasks, the language model’s advertising knowledge is enhanced, contributing to better performance in the next ad generation task.
@@ 4.3.2. Business Objectiveness Alignment
The primary and auxiliary tasks assist the LLM in recognizing user intent and interests, thereby generating ads that users are more likely to click on or convert. However, due to the LLM’s inherent lack of business context, the ads it generates may have limited business value, contributing minimally to cost and Gross Merchandise Volume (GMV). To address this, we introduce Direct Preference Optimization (DPO), which encourages the generation of ads with higher business value.
For a specific user $u$, the goal of DPO is to increase the likelihood of generating ads with high business value, denoted as $a_{h}$, while simultaneously reducing the probability of generating ads with lower business value, denoted as $a_{l}$. To construct the chosen pair $\langle u,a_{h}\rangle$ (preferred ad) and the reject pair $\langle u,a_{l}\rangle$ (less preferred ad) for DPO, we leverage the Expected Cost Per Mille (ECPM) metric during the ranking phase.
Specifically, for a given user $u$, we select two potential ads, $a_{1}$ and $a_{2}$, that are likely to result in a click or conversion. We then calculate the ECPM for each ad. The ad with the higher ECPM is considered the high-value ad $a_{h}$, while the ad with the lower ECPM is considered the low-value ad $a_{l}$. Through DPO, the LLM enhances its ability to recognize business value, thereby motivating the model to balance user intent with business objectives during ad generation.
Figure 3. Framework of Efficient System Deployment.
4.4. Efficient System Deployment
@@ 4.4.2. TensorRT LLM Acceleration
2) TensorRT-LLM Quantization LLM inference generally involves two stages: prefill and generate. Prefill is typically the computational bottleneck, while the generate stage is more memory-bound. However, in advertising scenarios with a large beam width and batching functionality supported by the inference engine, the generate stage also becomes a computational bottleneck. Therefore, we adopt quantization schemes that focus on activation values, such as smooth quantization (‘w8a8c8‘) and FP8 quantization (‘w8a8c8‘), to mitigate these bottlenecks.
@@ 5.2. Offline Performance
Variants HR@1 HR@4 HR@8 NDCG@4 NDCG@8
LEADRE 0.0764 0.1567 0.2021 0.1199 0.1360
w.o. content 0.0660 0.1343 0.1773 0.1029 0.1181
-13.60% -14.28% -12.26% -14.19% -13.15%
w.o. summary 0.0760 0.1543 0.2043 0.1181 0.1358
-0.62% -1.54% +1.08% -1.52% -0.18%
5.2.1. Effectiveness of Prompt Components
5.2.2. Effectiveness of Tuning Tasks and Their Order
@@ 5.2.3. Effectiveness of DPO
5.2.4. Effectiveness of S-IDs
To assess the effectiveness of the ads’ S-IDs, we compare the performance of different embedding models (emb. models). We first sampled 100 ads from each of the 48 ad categories and retrieved the Top-K nearest ads for each sampled ad based on embedding similarity. The accuracy is defined as the proportion of Top-K retrieved ads that belong to the same category as the original ad. The performance comparison is presented in Table 3. Our observations indicate that the E5 model demonstrates the best performance on this task, while the Hunyuan embedding achieves competitive results.
@@ 5.3. Online Performance
We conducted a 20% traffic A/B test on both the Tencent WeChat Channels and Tencent WeChat Moments display advertising systems to evaluate the effectiveness of LEADRE. Over several weeks’ A/B test on WeChat Channels, LEADRE demonstrated a 1.57% increase in Gross Merchandise Value (GMV) for serviced users. Similarly, several weeks’ A/B test on WeChat Moments showed a 1.17% increase in GMV for serviced users. These results highlight LEADRE’s capability to generate accurate, diverse, and high-value ads that contribute positively to business outcomes.

### v3

@@ 3.3. Trie-Tree Construction
We organize all ads using a trie-tree, where each path from the root to a leaf represents the S-IDs of an ad. This structure enables efficient ad organization and supports fast retrieval and prefix-based searches. The Trie-Tree Construction Algorithm (Algorithm 1) builds the trie by sequentially adding each S-ID from an ad as a child node, creating new nodes as needed, and marking the end of each ad’s S-ID sequence.
Example: Suppose we have the following ads, each represented by a sequence of S-IDs: Ad_66: $[a\_12,b\_7,c\_4]$; Ad_245: $[a\_12,b\_7,c\_14]$; Ad_112: $[a\_12,b\_6,c\_22]$. We construct the trie-tree as follows:
  1. (1)
    Initialize the trie with an empty root node.
  2. (2)
    Insert Ad_66: Add $a\_12$ as a child of root, then $b\_7$ under $a\_12$, and $c\_4$ under $b\_7$ (mark as ad end).
  3. (3)
    Insert Ad_245: $a\_12$ and $b\_7$ already exist; add $c\_14$ under $b\_7$ (mark as ad end).
  4. (4)
    Insert Ad_112: $a\_12$ exists; add $b\_6$ under $a\_12$, then $c\_22$ under $b\_6$ (mark as ad end).
The resulting trie-tree (Figure 1) allows for efficient ad retrieval and prefix-based searches, since ads sharing S-ID prefixes follow the same path from the root.
Algorithm 2 Constrained Decoding Using Trie-Tree
1:Trie-tree $T$ representing all ads, Beam search size $B$, S-IDs list length $(M+1)$, Current token probability distribution $P_{LLM}(c^{l}|\mathbf{c}^{1:l-1};\theta_{LLM})$ from the LLM.
2:Generated ad set of ads $\mathcal{B}$.
3:Initialize generated ad set $\mathcal{B}\leftarrow\{[h]\times B\}$, where $h$ is the head node of the trie-tree.
4:for $l=1$ to $M+1$ do
5:    Set current layer $cur\_layer$ as the $l$-th layer of trie-tree $T$.
6:    Initialize new generated set $\mathcal{B}^{\prime}\leftarrow\emptyset$.
7:    for each candidate list $\mathbf{c}\in\mathcal{B}$ do
8:         Retrieve valid next S-IDs from the children of node $c^{l-1}$ in $cur\_layer$: $\mathcal{C}_{valid}\leftarrow\text{children of}(c^{l-1})$.
9:         for each S-ID $c^{l}\in\mathcal{C}_{valid}$ do
10:             Obtain the score for S-ID $c^{i}$ from the LLM:
$P_{valid}(c^{l})\leftarrow P_{LLM}(c^{l}|\mathbf{c}^{1:l-1};\theta_{LLM}).$
11:             Append $c^{l}$ to the candidate list: $\mathbf{c}\leftarrow\mathbf{c}\cup c^{l}$.
12:             Add the updated list $(\mathbf{c},P_{valid}(c^{l}))$ to new set $\mathcal{B}^{\prime}$.
13:         end for
14:    end for
15:    Update the generated ad set: $\mathcal{B}\leftarrow\text{select\_top\_B}(\mathcal{B}^{\prime},B)$.
16:end for
17:return Generated ad set of ads $\mathcal{B}$.
@@ 3.4. Constrained Decoding
Unlike conventional LLM-based chatbots, generative retrieval in advertising requires generating a list of valid ads from a predefined set, ensuring both diversity and validity. To achieve this, we use constrained decoding with a trie-tree (Hokamp and Liu, 2017; Hu et al., 2019; Post and Vilar, 2018). This method ensures that generated ads are both likely according to the LLM and valid within the ad set. The Constrained Decoding Algorithm (Algorithm 2) expands candidate S-ID lists layer by layer through the trie-tree, selecting the top $B$ candidates at each step based on their scores, until all layers or the maximum length is reached.
Example: Suppose trie-tree is built from the following S-ID sequences: Ad_66: $[a\_12,b\_7,c\_4]$; Ad_245: $[a\_12,b\_7,c\_14]$; Ad_112: $[a\_12,b\_6,c\_22]$. Using constrained decoding with a beam size $B=2$:
  1. (1)
    Initialization: Start with an empty beam set at the root.
  2. (2)
    First Layer: Valid first S-ID from the trie-tree is $a\_12$ with $P(a\_12)=0.6$. Beam: $[a\_12]$ (score $0.6$).
  3. (3)
    Second Layer: For $[a\_12]$, valid next S-IDs are $b\_7$ and $b\_6$ with $P(b\_7)=0.5$, $P(b\_6)=0.4$:
    • $[a\_12,b\_7]$ (score $0.6\times 0.5=0.3$)
    • $[a\_12,b\_6]$ (score $0.6\times 0.4=0.24$)
    Keep both as $B=2$.
  4. (4)
    Third Layer:
    • For $[a\_12,b\_6]$: valid next S-ID is $c\_22$ with $P(c\_22)=0.8$, yielding $[a\_12,b\_6,c\_22]$ (score $0.24\times 0.8=0.192$).
    • For $[a\_12,b\_7]$: valid next S-IDs are $c\_4$ and $c\_14$ with $P(c\_4)=0.8$, $P(c\_14)=0.4$:
      • $[a\_12,b\_7,c\_4]$ (score $0.3\times 0.8=0.24$)
      • $[a\_12,b\_7,c\_14]$ (score $0.3\times 0.4=0.12$)
    After scoring, keep the top $B=2$: $[a\_12,b\_7,c\_4]$ and $[a\_12,b\_6,c\_22]$.
The algorithm returns the final beam set of S-ID sequences, ensuring that the generated ads are valid according to the trie-tree and have the highest predicted probabilities.
@@ 4.1. Overview
LEADRE effectively aligns user intent with business objectives, improving the generation of relevant and high-value ads in advertising environments. It’s framework consists of three novel modules:
(1) Intent-Aware Prompt Engineering: This module designs intent-aware ¡Prompt, Response¿ pairs that incorporate user profiles and behavior sequences, serving as a fine-tuning corpus for LLMs to learn user interests effectively.
(2) Advertising-Specific Knowledge Alignment: This module integrates auxiliary fine-tuning tasks and Direct Preference Optimization (DPO) to bridge the knowledge gap between LLMs and the advertising system, enhancing ad generation quality.
(3) Latency-Aware Model Deployment: This module combines both latency-tolerant and latency-sensitive service systems to ensure scalable and real-time performance during LLM deployment.
@@ 4.2.1. Prompt Components
To enable LLMs to capture user commercial intent effectively, we incorporate both long-term interests (derived from user profiles and historical behaviors) and short-term interests (based on recent behaviors) into prompt design. Short-term interests model dynamic, immediate preferences. For example, a user planning to purchase a phone would likely click phone or phone accessory ads within a brief period. However, such interests are inherently noisy and transient (e.g., after purchasing the phone, the user would cease clicking similar ads). Conversely, long-term interests capture stable, enduring preferences. For example, a student user would consistently engage with educational coaching ads over extended periods. Specifically, The prompt is assembled with the following key components:
@@
The following is an instruction describing a task. Please give a response to complete this request appropriately. 22-year-old male, resident in Haidian, Beijing, with a bachelor’s degree, working in Internet industry, with a medium consumption level. The categories that have been frequently interacted recently are (format: category^interaction times): emotion^115 times; entertainment^28 times; mental health^8 times; education^6 times; finance^6 times; The most recent interaction behavior sequence details (format: time^behavior type^title) are 32 days ago^play short video^reality; 31 days ago^click on ad^¡a_51, b_10, c_67, d_93, e_0¿; 27 days ago^play short video^shuttlecock; 25 days ago^Play short video^Emotional/psychological age test;22 days ago^Conversion ad^¡a_243, b_136, c_23, d_245, e_0¿;19 days ago^Click ad^¡a_164, b_243, c_38, d_88, e_0¿;16 days ago^Conversion ad^¡a_51, b_10, c_67, d_93, e_0¿. what ad will the user be interested in next time?
@@ 4.3.1. Semantic Alignment
Since LLM lacks prior knowledge of ads and their S-IDs, direct training on the main task encourages rote memorization rather than genuine understanding. This likes teaching calculus to students without foundational mathematics, which often results in solution memorization instead of conceptual comprehension. To address this, we introduce auxiliary tasks to guide LLMs learn the basic understand of advertising system and S-IDs of ads(Zheng et al., 2024). The auxiliary tasks consist of two components: an explicit alignment task and an implicit alignment task.
The explicit alignment task enables LLM to acquire S-ID knowledge by predicting S-IDs from detailed ad descriptions. This task directly aligns the language model with the ad system by making the LLMs understand how textual descriptions map to the corresponding ad S-IDs. A sample prompt-response pair for this task is shown below:
@@
In this example, LLM is provided with a detailed textual description of an ad, including the product’s name, type, brand, and category. The model’s task is to map this description to the corresponding S-IDs, which are internal identifiers used by the ad system. By adopting this task, LLM will have a better understanding about the S-IDs and ads’ feature, which plays a basic ability of generating ads.
The implicit alignment task develops advertising knowledge within user interaction contexts. This task builds on the Next Ad Generation mechanism described in the main task but replaces the S-IDs in the user’s ad interaction history with the ad descriptions. Instead of predicting the next ad based on a sequence of S-IDs (as in the main task), the model now predicts the next ad using textual descriptions of the ads. A sample ¡Prompt, Response¿ pair for this task is shown below:
[Prompt]: The following is an instruction describing a task. Please give a response to complete this request appropriately. $\cdots$ The most recent interaction behavior sequence details (format: time^behavior type^title) are 32 days ago^play short video^reality; 31 days ago^click on ad^Bali 7-Day Tour; 27 days ago^play short video^shuttlecock; 25 days ago^Play short video^Emotional/psychological age test;22 days ago^Conversion ad^psychological test;19 days ago^Click ad^Educational Counseling;16 days ago^Conversion ad^Tarot reading. What ad will the user be interested in next time? [Response]: ¡a_122, b_28, c_35, d_15, e_0¿
This subtle modification enhances the LLMs’ ability to associate ad descriptions with user behaviors, further aligning the model with the advertising system in a more implicit manner.
@@
Figure 3. Framework of Latency-Aware Model Deployment.
@@ 4.3.2. Business Objectiveness Alignment
Since LLMs inherently lack business context, they may treat all advertisements as having equivalent value. However, in advertising systems, each ad possesses distinct business value that directly impacts system revenue. When tuned exclusively using prior tasks, the generated ads often demonstrate limited commercial value, contributing minimally to both cost and Gross Merchandise Volume (GMV). To address this limitation, we employ Direct Preference Optimization (DPO) to encourage generating ads with higher business value.
Since LLMs inherently lack business context, they may treat all advertisements as having equivalent value. However, in advertising systems, each ad possesses distinct business value that directly impacts system revenue. When tuned exclusively using prior tasks, the generated ads often demonstrate limited commercial value, contributing minimally to both cost and Gross Merchandise Volume (GMV). To address this limitation, we employ Direct Preference Optimization (DPO) to encourage generating ads with higher business value.
As defined in Section 3.1, for a specific user $u$, we aim to predict the next interacted ad $a^{L+1}_{u}$ based on user behavior sequence $\mathcal{S}_{u}$. When considering two equally appealing ads $a_{1}$ and $a_{2}$ where $a_{1}$ holds greater business value, recommending $a_{1}$ is preferable due to its higher revenue potential. We therefore implement DPO to increase the probability of generating high-value ads ($a_{h}$) while reducing probability for lower-value ads ($a_{l}$). DPO preference pairs $\langle u,a_{h},a_{l}\rangle$ are constructed using the Expected Cost Per Mille (ECPM) metric from the ranking phase.
Specifically, for a given user $u$, we select two potential ads, $a_{1}$ and $a_{2}$, that are likely to result in a click or conversion. We then calculate the ECPM for each ad. The ad with the higher ECPM is considered the high-value ad $a_{h}$, while the ad with the lower ECPM is considered the low-value ad $a_{l}$. The training loss associated with DPO can be defined as:
$\mathcal{L}_{DPO}=-\mathbb{E}_{(u,a_{h},a_{l})\sim\mathcal{D}}\left[\log\sigma\left(\beta\frac{\pi_{\theta}(a_{h}|u)}{\pi_{ref}(a_{h}|u)}-\beta\frac{\pi_{\theta}(a_{l}|u)}{\pi_{ref}(a_{l}|u)}\right)\right]$
where $\mathcal{D}$ is the set of collected DPO training triplets, $\pi_{\theta}$ is the LLM tuned by DPO loss, $\pi_{ref}$ is the LLM tuned by primary and auxiliary tasks, $\sigma(\cdot)$ is the sigmoid function, and $\beta$ is a hyper-parameter. This loss function aims to maximize the likelihood of generating high-value ads while minimizing the likelihood of low-value ads, thereby aligning the LLM’s output more closely with business objectives.
4.4. Latency-Aware Model Deployment
@@ 4.4.2. TensorRT LLM Acceleration
2) TensorRT-LLM Quantization: LLM inference generally involves two stages: prefill and generate. Prefill is typically the computational bottleneck, while the generate stage is more memory-bound. However, in advertising scenarios with a large beam width and batching functionality supported by the inference engine, the generate stage also becomes a computational bottleneck. Therefore, we adopt quantization schemes that focus on activation values, such as smooth quantization (‘w8a8c8‘) and FP8 quantization (‘w8a8c8‘), to mitigate these bottlenecks.
@@ 5.1. Experiment Setup
Variants HR@1 HR@4 HR@8 NDCG@4 NDCG@8
LEADRE 0.0764 0.1567 0.2021 0.1199 0.1360
w.o. content 0.0660 0.1343 0.1773 0.1029 0.1181
-13.60% -14.28% -12.26% -14.19% -13.15%
w.o. summary 0.0760 0.1543 0.2043 0.1181 0.1358
-0.62% -1.54% +1.08% -1.52% -0.18%
@@ 5.2. Offline Performance
5.2.1. Effectiveness of Prompt Components
5.2.2. Effectiveness of Tuning Tasks and Their Order
@@ 5.2.3. Effectiveness of DPO
5.2.4. Effectiveness of Feature Embedding
To assess the effectiveness of the ads’ feature embedding used in ads indexing, we compare the performance of different embedding models (emb. models). We first sampled 100 ads from each of the 48 ad categories and retrieved the Top-K nearest ads for each sampled ad based on embedding similarity. The accuracy is defined as the proportion of Top-K retrieved ads that belong to the same category as the original ad. The performance comparison is presented in Table 3. Our observations indicate that the E5 model demonstrates the best performance on this task, while the Hunyuan embedding achieves competitive results.
@@ 5.2.4. Effectiveness of Feature Embedding
Table 5. Performance comparison of different codebook size and quantity.
[…]
Max collision
[…]
Average
[…]
3 1024 4.79% 44 1.0
512 7.95% 46 1.0
256 16.38% 67 1.0
4 1024 1.74% 25 1.0
256 3.34% 25 1.0
Table 6. Performance comparison of different S-Ids variants.
Variants HR@1 HR@4 HR@8 NDCG@4 NDCG@8
S-IDs 0.0800 0.1526 0.1975 0.1194 0.1353
Original Text 0.0306 0.0551 0.0679 0.0439 0.0487
-61.75% -63.89% -65.62% -63.23% -64.01%
Compressed Text 0.0572 0.1102 0.1317 0.0861 0.0967
-28.50% -27.79% -33.32% -27.89% -28.53%
5.2.5. Sensitivity and Effectiveness of S-IDs
[…]
Additionally, to assess the effectiveness of the S-IDs, we removed them and generated both the original and compressed text of ads using the LLM. The performance results are presented in Table 6. The findings indicate a significant performance drop when generating either the original or compressed text, highlighting the importance of S-IDs. Furthermore, the compressed text outperformed better than the original, suggesting that the original text contains redundancy and noise.
@@ 5.2.6. Effectiveness of User Behavior
5.2.7. Summary
In this section, we evaluated the impact of various components, including prompt engineering, tuning tasks, and S-IDs. Our findings are summarized as follows: (1) The S-IDs and the main tuning task are fundamental for ad generation; removing either results in the largest performance drops. (2) Auxiliary tasks, such as explicit and implicit alignment tasks and Direct Preference Optimization (DPO), contribute less than the S-IDs and main task, but still provide a crucial 9% performance increase. (3) User behaviors from the content and advertising domains are the most significant contributors to ad generation compared to other prompt components.
@@ 5.3. Online Performance
We conducted a 20% traffic A/B test on both the Tencent WeChat Channels and Tencent WeChat Moments display advertising systems to evaluate the effectiveness of LEADRE. LEADRE acts as a complementary retrieval sub-brunch. Over several weeks’ A/B test on WeChat Channels, LEADRE demonstrated a 1.57% increase in Gross Merchandise Value (GMV) for serviced users. Similarly, several weeks’ A/B test on WeChat Moments showed a 1.17% increase in GMV for serviced users. These results highlight LEADRE’s capability to generate accurate, diverse, and high-value ads that contribute positively to business outcomes.
LEADRE operates as a retrieval sub-branch, collaborating with other LTR branches to complete the retrieval process. To highlight the competition between branches, we introduce the consumption. Consumption is the proportion of exposed ads exclusively retrieved by a specific branch relative to all exposed ads. Online A/B test shows that LEADRE achieves a consumption ratio of 7.33% (Wechat Channels). This demonstrates its ability to retrieve ads missed by other branches, directly contributing to the increase in GMV.